OpenStudy (anonymous):

A gamma ray with an energy of 3.40 × 10-14 joules strikes a photographic plate. We know that Planck's constant is 6.63 × 10-34 joule·seconds. What is the frequency of the photon? Select one of the options below as your answer: A. 5.13 × 1019 hertz B. 5.60 × 1017 hertz C. 2.45 × 1017 hertz D. 2.45 × 10-12 hertz E. 7.80 × 10-19 hertz

OpenStudy (anonymous):

can someone plug this in for me and ill solve it im bad at doing that stuff

OpenStudy (anonymous):

The relation is (which u must be knowing) \(E=h\nu\) where \(E \) is energy which is given to be as 3.40×10\(^{-14}\) J. \(h\) =6.63×10\(^{-34}\) Js. Just put these values in the above and find \(\nu\),the frequency.

OpenStudy (anonymous):

ok ty ill try it out now

OpenStudy (anonymous):

what would be the v?

OpenStudy (anonymous):

You have to find it. It is the frequency

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

should I use a regular calcuklator or could it work on mathway.com?

OpenStudy (souvik):

use a scientific calculator

OpenStudy (anonymous):

ok ill try it out

OpenStudy (anonymous):

on the scientific calculator on the computer how do I do 10^-14?

OpenStudy (anonymous):

-14 on top

OpenStudy (souvik):

do not calculate the powers of 10...you can make out without calculator

OpenStudy (anonymous):

idk if i can tbh

OpenStudy (anonymous):

I really didnt learn this...

OpenStudy (souvik):

3.40/6.63=?

OpenStudy (anonymous):

its A?

OpenStudy (souvik):

what did u get?0.513?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

0.5128

OpenStudy (anonymous):

well u just round

OpenStudy (souvik):

now 10^-14/10^-34=10^20 so the answer is A

OpenStudy (anonymous):

ok cool ty